OpenStudy (anonymous):

A car travels up a hill at the constant speed of 40 km/h and returns down the hill at the speed of 60 km/h. Calculate the average speed for the round trip.

OpenStudy (shamim):

i think the result will b 50km/h

OpenStudy (anonymous):

Yes becuase 60+40=100, then 100/2 is 50

OpenStudy (shamim):

ya i think so

OpenStudy (anonymous):

yes no doubts!!! your answer is 50!!!!

OpenStudy (anonymous):

lets say that the way you travel is 60 km (since distance isn't given, it can be pretty much any you would like to take, best is you make it S), then you get that you travel for 1.5 hours towards +1 hour backwards. so you get it like : S=V*t V= S/t (the traveled distance can be described as S= 40*1.5 +60*1 ) V= (4*1.5 +60*1)/2.5 =120/2.5=48 km/h (this is example with traveled distance to one side=60) So, the answer is this formula: V(average)= (60*t1+40*t2)/(t1+t2) Or i might be mixing up everything, not sure on that ( 50 is the speed of car without gravity's acceleration from my point of view).